OpenStudy (anonymous):

The operation* is defined as: x*y= (x^2 +y^2+2xy)/(x+y). If x*8=3, then which of the following is a possible value of x? -11, 8, 11, -5, or 3

OpenStudy (anonymous):

first note that (x^2 +y^2+2xy) = (x+y)^2 then u get: x*y= x+y [1] given that when y=8, x*y=3 therefore: equation [1] becomes: 3=x+8 this implies that x=-5